OpenStudy (anonymous):

Write in point slope form Through: (-1, -5), slope=5

OpenStudy (gorv):

y-y1=m(x-x1) y-(-5)=5(x-(-1))

OpenStudy (gorv):

y+5=5(x+1) y+5=5x+5